OpenStudy (anonymous):

Rewrite in simplest radical form 1/x^-3/6.

OpenStudy (anonymous):

\[\frac{ 1 }{ x^{\frac{ -3 }{ 6 }} }\]

OpenStudy (welshfella):

Hint: 1 / a^-2 = a^2

OpenStudy (anonymous):

So, x^3/6? Do you just make it positive?

OpenStudy (welshfella):

yes that is correct now simplify 3/6

OpenStudy (welshfella):

3/6 = ?

OpenStudy (anonymous):

How do you simplify that? ?/3?

OpenStudy (welshfella):

3/6 = 1/2

OpenStudy (anonymous):

May you explain how you did that, for me.

OpenStudy (welshfella):

its just making the fraction simpler by dividing top and bottom by a common factor The common factor in this case is 3. so